Find any numbers for which the rational expression is undefined:
v-2/5v-9

Answer by stanbon(75887)

Find any numbers for which the rational expression is undefined:
(v-2)/(5v-9)
------
The denomiator cannot be zero.
Solve 5v-9 = 0
5v = 9
v = 9/5
------
The expression is undefined for v = 9/5
------------------
Cheers,
Stan H.